# Are Generative AI systems Capable of Supporting Information Needs of Patients?

### v1

@@ 4.2. Inductive Thematic Analysis
4.3. Frequency of Themes across Radiologist-Patient Interactions
Figure 3 describes the proportion of interview content each theme occupied in each participant’s interaction with the radiologist. Although there are variations in how much each participant’s conversation incorporates each theme, there appear to be several consistencies across participants. The following are a few noteworthy observations from the graph:
@@ 4.3. Frequency of Themes across Radiologist-Patient Interactions
Figure 4. Stacked bar chart describing the proportion of each theme in each participant’s interaction with the radiologist, with the portion of the conversation occupied by each theme on the y-axis, and participants along the x-axis.
5. Generative AI Evaluation
To effectively use Generative AI systems to support patients’ understanding of their medical reports, they must be able to respond reliably, factually, and meaningfully to diverse questions asked by patients. Here, we evaluate if the current generative AI models can meet this desiderata. Our evaluation focuses on two state-of-art foundational models for the medical visual questions answering task: MedFlamingo (moor2023med, ) and ChatGPT-4V  (yang2023dawn, ), elaborated in Section 2.
5.1. Method
Our evaluation paradigm is based on the materials in our study (Section 4) and our inductive thematic analysis of the patient-provider conversation. From the $10$ themes identified in Section 4.2, we selected $4$ themes to use in our evaluation of generative models above: theme 4, selecting a treatment plan; theme 3: locating issues mentioned in the radiology report to the CT scan; theme 5, understanding medical terminology from radiology reports, and theme 6, diagnostic follow-up plan. We chose these $4$ themes because they allowed for the clearest formulation of meaningful question-answer pairs. The remaining themes were largely observed in parts of the radiologist’s explanations to answers to questions from the selected $4$ themes. From each of these themes, we selected $5$ questions that were asked by the participants representative of the diversity in each theme, maximizing variety both in terms of the content of the question and the participant asking the question. These questions and the corresponding answers given by the radiologist formed the evaluation set for the generative AI models.
Figure 5. An example of a medical visual question answering problem in our evaluation
We structured the evaluation as a visual questions-answering problem shown in Figure 5. A problem consisted of (1) two images from radiology scans, (2) corresponding radiology report, and (3) a question. A major concern with the models we evaluated is the character limit on the length of the prompt - if the prompt becomes too lengthy, it is truncated internally in the model. To fit the medical report in the given character limit, we manually summarized it and validated it for accuracy and correctness. Another challenge is how patients phrase their questions in a natural conversation. Often, the questions are not grammatical, have disfluencies, and may be convoluted. Generative AI models we evaluated are not designed to handle such natural conversational patters. Therefore, we paraphrased the question such that they are grammatical and clearly stated. Our final evaluation set consisted of $20$ question-answer problems. Each model was prompted with a problem and the answer it generated was recorded.
It is noteworthy that generative AI models are malleable and their behavior can be modulated through in-context learning and few-shot prompting such that it follows a desired style, format, or reasoning approach. Our evaluation protocol doesn’t leverage this capability, but evaluates the nascent behavior of these models. However, it has the advantage of being simple and straightforward leading to a clearer analysis approach. In future, we will study various in-context and few-shot learning prompts and their impact on the model responses in a principled fashion.
5.2. Analysis
We analyzed the answers generated by MedFlamingo and ChatGPT-4V with respect to those provided by the radiologists during our interview study on two dimensions, qualitatively. Figure 6 shows an example question from our evaluation set and corresponding answer provided by the expert as well as those generated by MedFlamingo and ChaptGPT-4V.
Figure 6. Question and corresponding answers provided by the expert, MedFlamingo, and ChatGPT-4
Summary Statistics
In our evaluation set, the radiologist’s responses had $M=136.95$ ($SD=75.53$) words. MedFlamingo’s responses tended to be significantly more concise with $M=52.05$ ($SD=22.28$) words. ChatGPT4 was the most verbose of all responses with $M=301.42$ ($SD=78.48$) words on average.
Evaluation Criterion
We qualitatively evaluated the answers generated by Generative AI models on two dimensions: correctness and relevance. Correctness evaluates if the answers generated by a model was judged to be right given the patient’s by an expert evaluator (a radiologist and a co-author on this paper). We adopt a particularly strict criterion for correctness - whenever the expert identified even a single non-factual sentence in the response, it was marked to be incorrect. Such strict evaluation is necessary for critical applications of generative AI models that have been known to hallucinate information which may be dangerous in our usecase. Relevance evaluates how well the question is answered. For this, we characterized the information in sentences in the answers as: direct response, sentences that contain a clear answer to the question asked; relevant elaboration, sentences that elaborate the answer by providing additional, supporting information; irrelevant elaboration, sentences that contain information irrelevant to the question asked; and finally, superfluous, sentences with little information about the question. Figure 6 shows an example of such characterization where text in blue indicates direct response, green - relevant elaboration, orange - irrelevant elaboration, and purple - superfluous text. The characterization was done by authors who are not medical experts and appropriate evaluators of the relevant dimension.
@@ 5.2. Analysis
On the correctness dimension, we found that out of $20$ questions, MedFlamingo was judged to be incorrect on $14$ ($70\%$ error rate) while ChatGPT-4V on $7$ ($35\%$ error rate). Errors manifested due to a variety of causes including: inablility to relate causal factors to diseases, incorrectly identifying risks and consequences, inability to distinguish generic textbook information from specifics of the case, and factual incorrectness due to hallucinations. In one case, ChatGPT identified itself as a radiologist. In at least one instance where both MedFlamingo and ChatGPT were judged to be correct, the expert preferred the answer generated by MedFlamingo because it was concise. In a couple of cases where ChatGPT4’s answer was judged to be correct, the expert noted that it “it rambles on and on about risk stratification and coronary calcification when the patient simply asked where the nodule is located”. In another case, the highly detailed answer provided by ChatGPT-4V prompted the expert to say “but given the detailed answer it gave, I felt that it would be necessary for it to have mentioned what staging system (such as TNM staging) that it used”. Both these cases indicate that ChatGPT-4V is unable to distinguish between relevant and irrelevant information and indiscriminately produces both.
@@
Our analysis of the relevance dimension revealed several insights about how doctors address their patient’s questions. In a large number of cases ($11$) the doctor’s response consisted of sentences that directly answer the patient question followed by a relevant elaboration. An example is shown in Figure 6 where after explaining that the majority of nodules do not turn out to be cancer, the doctors describes the criteria that are used to evaluate nodules. Other elaborations included demonstrations, impact of the disease the patient can expect to see, what further evidence may manifest, how a diagnosis may be made, and how a treatment plan may be developed. This finding hints at the mixed-initiative nature of the patient-provider dialog where the role of the provider is not just to answer questions but also educate the patient so that they feel empowered to take informed decisions about their health. We didn’t observe any other information categories in the radiologist’s response.
MedFlamingo always provided a direct response to the question with no elaborations. Its answers tended to be very concise and expressed in clinical language. Additionally, in $6$ responses it used first-person pronouns to recommend a specific action. For example, in response to the questions “Will there be a follow-up in 3 months, or will you contact me again if there is an acute need for treatment?” ”I would recommend a follow-up CT scan in 3 months. If there is any acute treatment need, we will contact you.”. This is in stark contrast with the radiologist’s response who doesn’t express their medical opinion but describes what potential next steps might be.
ChatGPT-4V generated long answers with elaborations. An example is shown in Figure 6 where the relevant elaborations are shown in green and the irrelevant ones in orange. Explanation of how nodule’s size or appearance influences the diagnosis of cancer is relevant for the patient. However, elaborations about emphysema and coronary artery calicification are not pertinent to the question asked, even though they were mentioned in the report. Additionally, it generated some superfluous information shown in purple that prefaces its actual response but doesn’t add any useful information to the response. We found that it was unable to distinguish relevant elaborations from irrelevant information that it generated based on terms used in the report.
@@
Both AI systems struggled with responding to questions in theme 3: locating issues mentioned in the radiology report to the CT scan. An illustrative example is shown in Figure 7. To answer the question, the radiologist responded by first describing how to find relevant areas - nodules - in the scan and then, a demonstrative elaboration highlighting those in the scan to ground the caregiver’s understanding. The demonstrative elaboration was copiously missing in both MedFlamingo and ChatGPT-4V responses. It is to be expected. Medical visual question answering systems are not designed to highlight which piece of visual evidence contributed to answer generation. Additionally, MedFlamingo introduced clinical language and attempted to generate a clinically comprehensive answer about the nodule’s location which is unlikely to support the caregiver’s understanding. ChatGPT-4V on the other hand generated generic information about how to find various types of nodules in the scans. Both missed the fact that there are in fact two nodules in the scan as is clear from the radiologist’s response.
@@
Figure 7. Theme 3 question and corresponding answers provided by the expert, MedFlamingo, and ChatGPT-4V
Finding 4
We also found that both AI systems were sensitive to the way the question was framed and demonstrated an over-reliance on the textual report as opposed to applying reasoning to interpret the images. For example, when asked the question ”How do I find a nodule here?”, in addition to providing a brief description of how to spot a nodule in a scan, both models went on to describe only one of the two nodules present. The nodule they described was the one that was associated with the phrase ”solid nodule” in the report, and the one they missed was the one described as ”ground glass nodular consolidation”, suggesting that they simply looked to the terminology in the report instead of studying the image. When asked ”Can we discuss the specific areas where the suspicious nodules might be located?”, both models mention both the nodules. However, they struggled to describe the location of the ”ground glass nodular consolidation” as its location was not explicitly stated in the report.”
In general, we found that MedFlamingo tended to generate concise answers in clinical language which may be hard for the patients and caregivers to understand. On the other hand, ChatGPT-4V tended to generate long answers with generic descriptions of terms used in the report, generic lists of risks, treatment options, etc. without any focus on the specifics of the case under discussion and with no ability to filter relevant details from irrelevant ones.
6. Discussion, Conclusions, and Future Work
@@
In this paper, we study if current state-of-art in generative AI systems can support patients (and caregivers) as they attempt to understand their scans and reports. To structure our evaluation, we conducted a need-finding study of caregiver-radiologist interactions. We identified $10$ themes of diverse information types that such interactions are typically comprised of. We used $4$ themes to create an evaluation question-answer dataset and evaluated two state-of-art models: ChatGPT-4V and MedFlamingo on two dimensions relevant to our usecase, correctness and relevance.
@@
Our findings indicate that despite significant optimism about generative models in medical contexts, the current state-of-art doesn’t sufficiently address patients’ information needs. Surprisingly, we found that these models have high error rates ($35\%$ for ChatGPT-4V, $70\%$ in MedFlamingo) in our evaluation set built on question asked in real interactions. This finding is concerning - patients are not experts of medicine and consequently, cannot identify misinformation in AI-generated responses. Any system that is deployed must guaranteed to provide factual information. The models were also deficient as information systems and frequently produced irrelevant and superfluous elaborations that didn’t specifically answer the question asked but rather elaborated terms used in the medical report. This finding suggests that it is unlikely that patients will be able to use this system to access useful information.
@@
Our findings also highlight a critical gap in the evaluation paradigm prominent in AI & ML and the challenges of accountable deployment of generative AI systems. In machine learning literature (bubeck2023sparks, ), these models are shown to have high accuracy on curated, task-agnostic datasets and claimed to have human-level performance. However, to be at par with human experts, the models not only have to answer questions but also operationalize this knowledge in practical, real-world tasks such as helping patients understand their illness and determine a treatment plan - a task doctors and healthcare worker routinely perform. Our work paves the way for task-specific evaluation protocols that can spur the generative AI movement in a beneficial direction.
@@
There are several avenues for future work. First, we will expand the study described here to cover additional medical contexts to come up with an exhaustive set of informational needs patients have. We will use this set to guide the development of an evaluation benchmark that reflects the structure of this real-world task. Second, prominent metrics such like BLEU (papineni2002bleu, ), ROGUE (lin2004rouge, ), BERT Score (zhang2020bertscore, ) while commonly used to evaluate LLM responses, fall short of measuring if the responses are useful in addressing human needs. To address this shortcoming in AI & ML evaluation paradigm, we will develop a quantitative measurement scale based on our the qualitative evaluation paradigm proposed here. Finally, we are invested in developing intelligent assistive systems for the healthcare. Our future work will explore how the shortcoming highlighted in this paper can be alleviated with modern and classical AI methods to develop accountable and transparent systems.
Research Ethics and Social Impact
Research materials used for conducting the study in this paper are based upon sensitive medical data obtained from real patients accessing the public university hospital for care. The data was modified and redacted as per the hospital’s guidelines before it was shared with the research team. This step is critical for protecting patients’ privacy.
The study described in this paper is sensitive in nature. It asks the participants to imagine themselves in a role of helping a close relative with medical decisions pertaining to a serious illness. This simulation can trigger panic and anxiety from past trauma if the participant or their loved ones have gone through similar experiences. Our recruitment emails highlighted this characteristics of the study as a trigger warning. Our consent forms also highlighted this concern, encouraging potential participants to reflect on their experiences and only consent if they felt comfortable. Finally, our consent form informed the participants that they can terminate the study at any point without incurring any adverse consequences. The IRB at the research company approved this study with an expedited review process and classified it as minimal risk.
The authors on this paper belong to diverse intellectual backgrounds; from AI & ML, to cognitive science, HCI, and education, to medicine. Our joint interest in building intelligent systems can lead to an optimism bias about the usefulness of AI & ML in healthcare. To overcome that bias, in this paper, we apply a critical lens to evaluate the use of this technology in the healthcare context. Our analysis is centered on human needs and evaluates the state of art on its suitability in addressing those needs.
Our work is best understood as a proposal for an evaluation approach for generative AI systems. An adverse impact this paper can have on the research community is accepting this approach as exhaustive, complete, and the only correct way. The research on evaluating generative AI system is nascent and rapidly evolving. Diverging perspectives on how to measure system performance are useful for the advancement of these systems.

### v2

@@ 4.2 Inductive Thematic Analysis
4.3 Frequency of Themes across Radiologist-Patient Interactions
Figure 4 depicts the proportion of various themes in each participant’s interaction with the radiologist. Although there are variations in how much each participant’s conversation incorporates each theme, there appear to be several consistencies across participants. The following are a few noteworthy observations:
@@ 4.3 Frequency of Themes across Radiologist-Patient Interactions
Figure 4: Stacked bar chart describing the proportion of each theme in each participant’s interaction with the radiologist, with the portion of the conversation occupied by each theme on the y-axis, and participants along the x-axis.
5 Generative AI Evaluation
Successful adoption of generative AI systems in patient-facing applications requires that they must respond reliably, factually, and meaningfully to questions asked by patients. In this section, we evaluate two state-of-art, commercially-deployed generative AI (GenAI) systems : Claude 3.5 Sonnet [19] and ChatGPT 4o [48]. Both of these systems are built upon language models that are top-scores in standardized multi-task understanding benchmarks such as MMLU-Pro [49]. MMLU-Pro is a question-answering dataset containing $12032$ questions and multiple choice answers from health, psychology, and biology domains in addition to other fields. Our goal is to explore if these highly successful models can address questions humans ask naturally as they perform a sensemaking task.
5.1 Study
5.1.1 Conversational MultiModal Dataset
We, first, built a small multimodal dataset of question-answer pairs from the formative interview study and corresponding thematic analysis 4. Our dataset represents $5$ themes;
  1. 1.
    Theme 3: locating issues mentioned in the radiology report to the CT scan,
  2. 2.
    Theme 4: selecting a treatment plan,
  3. 3.
    Theme 5: understanding medical terminology from radiology reports,
  4. 4.
    Theme 6: diagnostic follow-up plan, and
  5. 5.
    Theme 7: risk factors and lifestyle changes.
We chose these $5$ themes because they allowed for the clearest formulation of meaningful question-answer pairs. The remaining themes were largely observed in parts of the radiologist’s explanations to answers to questions from the selected themes. For each theme, we selected $5$ questions that were representative of the diversity in the questions study participants asked. We paired each question with the long-form answer that the radiologist provided during the interview study.
We structured the evaluation dataset as a set of $25$ visual question-answering problems. Each problem consisted of (1) two images from radiology scans (in Figure 2), (2) corresponding radiology report capturing expert impressions (in Figure 2), and (3) a question asked by the patient. This setup requires the GenAI system under study to perform visual reasoning on the CT scan images and integrate that evidence with analytical understanding of the report to generate long-form answers.
5.1.2 Protocol
We accessed both Claude-3.5-Sonnet and ChatGPT-4o through their web interfaces. Generative AI systems are malleable and their behavior can be modulated through system instructions and various prompting techniques. Several of these prompting techniques (few-shot prompting, chain-of-thought reasoning) require the human to have expertise in interacting with generative AI systems and an understanding of their behavior. Patients are unlikely to have this expertise. Consequently, we studied two configuration that are most natural in our setting. In the first configuration (/config1), the system is provided with all background material (two scans and the report) and asked a question. This configuration reflects how a patient might use the publicly available GenAI systems. In the second configuration (/config2), we added a system prompt to guide AI’s behavior. The system prompt provides the context in which the question is asked and includes who is asking the question (who: a patient’s caregiver), what the questioner is looking for (why: they want to understand the scans and reports), and how the response should be generated (how: simple, non-clinical language, and a succinct answer). The configuration instructs GenAI systems that the questioner doesn’t have a background in clinical practice, biasing them to use colloquial language and descriptions. This configuration reflects the guidance a GenAI system might receive from a healthcare system administrator as the system is deployed.
Both GenAI systems were tested under the two configurations on all $25$ evaluation questions and their answers were recorded. After the system answered a question, the conversation history was reset. This was done to ensure independence in response generation. GenAI systems are autoregressive and generate answers by sampling from a token distribution based on which tokens were generated previously in the conversation. Resetting the history ensures that incorrect information in earlier response do not bias response generation for the question under study.
We complied the questions, expert answers, and responses generated by both configurations of the two GenAI systems understudy in an evaluation dataset [^1] Raw dataset is available at https://docs.google.com/spreadsheets/d/e/2PACX-1vTZB2sw1JadKKc1TXF8dzF-WYL7cOFta6E61_IIdyVvR1ncDzm0nceFpA3JFPwZ5acWVjDM8hpYB9Ki/pubhtml.
5.2 Analysis
5.2.1 Empirical Findings
Table 1 shows the average length of expert responses as compared to those produced by the GenAI systems under both experimental conditions. Natively, both Claude-3.5-sonnet and ChatGPT-4o tend to be verbose and produce significantly longer responses. However, when they are configured with the system prompt encouraging them to produce answers that are succinct, the response length reduces reliably for both.
Table 1: A table noting the average length of responses from a radiologist and variations of GenAI models: Claude and ChatGPT
Response type Avg. words (std. dev.)
Radiologist 146.2 (82.68)
Claude/config1 281.0 (35.85)
Claude/config2 167.88 (79.22)
ChatGPT/config1 178.36 (53.42)
ChatGPT/config2 134.12 (45.77)
Next, we evaluate if the GenAI system responses are similar to how a radiologist responds. Evaluating similarity of free-form text is a challenge in itself. Classical measures such as ROUGE score [50] rely on overlap in the words used in two pieces of text. However, difference in stylistic choices may result in lower scores even when the two texts have similar meaning.
To compare meaning-level similarity between expert and GenAI responses, we use the universal sentence encoder (USE) that encodes text into high dimensional vectors [51]. The model is trained and optimized for greater-than-word length text, such as sentences, phrases or short paragraphs making it suitable for our purposes. It is trained on a variety of data sources and a variety of tasks with the aim of dynamically accommodating a wide variety of natural language understanding tasks. The input is variable length English text and the output is a $512$ dimensional vector. Once the responses are converted into high-dimensional vectors, their cosine similarity ranges from 00 to $1$ indicating how close they are in the semantic space, $1$ when the responses are the same. In other words, the cosine similarity indicates the degree to which a GenAI system’s response is aligned with an expert.
Figure 5 shows how similar the responses generated by the experimental configurations of GenAI systems are to those of a human expert. The scale goes from dark blue to dark red (dissimilar to similar). We make the following observations.
  1. 1.
    There is a large variability in measured similarity. This variability indicates that generative AI systems are not in full alignment with the expert radiologist. In a majority of the cases, their responses are different from that of a radiologist scoring less than $0.75$ on cosine similarity.
  2. 2.
    Response similarity varies between the themes. Highest similarity are observed in Theme 5: Understanding Medical Terminology in the Radiology Report. This is not surprising given that clinical textbooks are included in the datasets LLMs are trained over. Lowest similarity responses are observed in Theme 3: Locating Issues Mentioned in Radiology Report to the CT scan. We investigate the reason for this observation in our qualitative analysis (4.2).
  3. 3.
    Configuration 2 that encourages the system to produce short text accessible to non-clinicians appears to reduce the similarity measure. Mostly, the impact is small. However in some cases (e.g, q3, q14 Claude; q21, q22 ChatGPT) it is especially pronounced. Reduction in similarity indicates that GenAI system do not know how to focus the generative process on relevant information. They can remove information that the radiologist considers important to communicate.
Figure 5: A heatmap plotting the semantic similarity of responses generated by Claude and ChatGPT under two experimental configurations to an expert’s response.
5.2.2 Qualitative Investigation
Similarity scores reported earlier indicate that there are differences between how a radiologist responds to patient queries and how a GenAI system approaches answer generation. However, the scores themselves don’t explain why or how these answers are different. In this section, we qualitatively analyze the difference between the responses provided by the radiologist from those generated by the GenAI systems.
Drawing inspiration from the Griean maxims [52] of effective conversation, we identify the following dimensions of evaluation.
  1. 1.
    Correctness follows the Gricean maxim of quality, where one tries to be truthful, and does not give information that is false or that is not supported by evidence. We engaged an expert evaluator (a radiologist and a co-author on this paper) to judge if the responses produced by the GenAI systems were correct. We adopted a particularly strict criterion for correctness - whenever the expert identified even a single non-factual sentence in the response, it was marked to be incorrect. Such strict evaluation is necessary for critical applications of generative AI models that have been known to hallucinate information which may be dangerous in our usecase.
  2. 2.
    Relevance and informativeness follow the Gricean maxim of relation (where one tries to be relevant, and says things that are pertinent to the discussion) and quantity (where one tries to be as informative as one possibly can, and gives as much information as is needed, and no more). We characterized each sentence in the answers as: direct response, sentences that contain a clear answer to the question asked; relevant elaboration, sentences that elaborate the answer by providing additional, supporting information; irrelevant elaboration, sentences that contain information irrelevant to the question asked; and finally, superfluous, sentences with little information about the question. Figure 6 shows an example of such characterization where text in blue indicates direct response, green - relevant elaboration, orange - irrelevant elaboration, and purple - superfluous text. The characterization was done by authors who are not medical experts and appropriate judges to evaluate if the answers are relevant and informative.
Figure 6: Theme 5 question (id=q1) and responses by the expert and GenAI systems. Blue denotes sentences that contain a direct response to the question; green, relevant elaborations; orange, irrelevant elaboration; and purple, superfluous text.
@@ 5.2.2 Qualitative Investigation
On the correctness dimension, we found that out of $25$ questions, Claude/config1 was judged to be incorrect on $10$ ($40\%$ error rate) while ChatGPT/config1 on $5$ ($20\%$ error rate) leading the expert to reject their answers. In two instances the expert accepted the responses while noting that the response didn’t answer the question but was correct. Errors manifested due to a variety of causes including: inablility to relate causal factors to diseases, incorrectly identifying risks and consequences, inability to distinguish generic textbook information from specifics of the case, and factual incorrectness due to hallucinations. The expert noted that the GenAI systems ramble on and often, include information that is not directly relevant to the question.
@@
Our analysis of the relevance & informativeness dimension revealed several insights about how doctors address their patient’s questions. The radiologist’s responses consisted of sentences that directly answer the patient question followed by a relevant elaboration. An example is shown in Figure 6 where after explaining that the majority of nodules do not turn out to be cancer, the doctors describes the criteria that are used to evaluate nodules. Other elaborations included demonstrations, impact of the disease the patient can expect to see, what further evidence may manifest, how a diagnosis may be made, and how a treatment plan may be developed. This finding hints at the mixed-initiative nature of the patient-provider dialog where the role of the provider is not just to answer questions but also educate the patient so that they feel empowered to take informed decisions about their health. We didn’t observe any other information categories in the radiologist’s response.
Both Claude-3.5-Sonnet and ChatGPT-4o produced long answers with elaborations. An example is shown in Figure 6 where the relevant elaborations are shown in green and the irrelevant ones in orange. Explanation of how nodule’s size or appearance influences the diagnosis of cancer is relevant for the patient. However, next steps are not pertinent to the question asked. Additionally, both generated some superfluous information shown in purple preface or summarize its actual response but doesn’t add any useful information to the response. Production of irrelevant elaborations and superfluous information directly violates of desiderata of relevance & informativeness and the Gricean maxims of relation and quantity.
@@
Both GenAI systems struggled with responding to questions in theme 3: locating issues mentioned in the radiology report to the CT scan, explaining the low similarity scores in Figure 5. An illustrative example is shown in Figure 7. To answer the question, the radiologist responded by first describing how to find relevant areas - nodules - in the scan and then, a demonstrative elaboration highlighting those in the scan to ground the caregiver’s understanding.
The demonstrative elaboration was copiously missing in both Claude/config1 and ChatGPT/config1 responses. They both generated generic guidance on how to locate the nodule in the scan and regurgitated language about nodules from the scan. It is to be expected. GenAI systems are not designed to guide the questioner’s attention to visual evidence contributed to answer generation. Further, Claude/config1 and Claude/config2 incorrectly believe that the nodule cannot be localized in the scans. Our observations call into critique the assumed ’multi-modality’ of GenAI system when these systems only consume visual information to guide response generation but cannot use the visual medium to communicate information.
ChatGPT/config2 generated an answer that is correct however, is entirely useless to the patient. It present banal/obvious information about how to localize the nodule. This finding indicates that we should revise our relevance and informativeness criteria to include banality.
@@
Figure 7: Theme 3 question (id=q7) and corresponding answers provided by the radiologist and GenAI systems
Finding 4 In several responses, the radiologist went beyond the exact question being asked and answered a higher level concern. This approach was particularly visible in Theme 6: diagnostic follow-up plan. An example is shown in Figure 8 where the patient enquires about expected expenses for various treatment options. The radiologist’s response not only compares different treatment options in terms of how expensive they are, it also provides them with other tradeoffs that the patient didn’t explicitly inquire about. Examples include observations such as radiation oncology requires the patient to come back several times and chemotherapy has side-effects that might require hospitalization. In contrast, both Claude-3.5-Sonnet and ChatGPT-4o focused on costs associated with treatment options, often elaborating to include less common or new therapies. In general we found the radiologist’s responses were geared towards helping the patients make decisions whereas GenAI systems’ responses were geared towards extensive enumeration of possibilities and definitions.
Figure 8: Theme 6 question (id=q15) and corresponding answers provided by the radiologist and GenAI systems
Finding 5 A meta-finding of our research is that the USE-based similarity metric isn’t sufficiently selective for our tasks. Consider the example in Figure 7 where it is easy to see that the radiologist’s responses in grounded in the CT scans and discusses localization of nodules. GenAI systems on the other hand produce responses containing different information. However, their similarity scores (Figure 5, line q7) are almost $0.7$ indicating a good degree match. This observation indicates that USE supports a topic-level similarity but cannot compare the content as a human would, despite being one of the standard measurement techniques in AI.
6 Discussion, Conclusions, and Future Work
@@
In this paper, we study if current state-of-art in generative AI systems can support patients (and caregivers) as they attempt to understand their scans and reports. To structure our evaluation, we conducted a need-finding study of caregiver-radiologist interactions. We identified $10$ themes of diverse information types that such interactions are typically comprised of. We used $4$ themes to create an evaluation question-answer dataset and evaluated two state-of-art models: Claude-3.5-Sonnet and ChaptGPT-4o both quantitatively, by measuring response similarity with a human expert and qualitatively, by judging the responses on Gricean maxims of effective communication.
@@
Our findings indicate that despite significant optimism about generative models in medical contexts, the current state-of-art doesn’t sufficiently address patients’ information needs. Surprisingly, we found that these models have high error rates ($20\%$ for ChatGPT-4o, $40\%$ in Claude-3.5-Sonnet) in our evaluation set built on question asked in real interactions. This finding is concerning - patients are not experts of medicine and consequently, cannot identify misinformation in AI-generated responses. Any system that is deployed must guaranteed to provide factual information. The models were also deficient as conversational information systems and frequently produced irrelevant and superfluous elaborations violating Gricean maxims. This finding suggests that it is unlikely that patients will be able to use this system to access useful information.
@@
Our findings also highlight a critical gap in the evaluation paradigm prominent in AI & ML and the challenges of accountable deployment of generative AI systems. In machine learning literature [53], these models are shown to have high accuracy on curated, task-agnostic datasets and claimed to have human-level performance. However, to be at par with human experts, the models not only have to answer questions but also operationalize this knowledge in practical, real-world tasks such as helping patients understand their illness and determine a treatment plan - a task doctors and healthcare worker routinely perform. Our work paves the way for task-specific evaluation protocols that can spur the generative AI movement in a beneficial direction.
There are several avenues for future work. The study described here is small (only $5$ participants) and limited to a single case of radiological imaging. Consequently, the themes identified here do not represent an exhaustive set of informational needs patients have. In doing this arguably limited study, we formulated a concrete approach of evaluating GenAI systems. Building upon our learnings, we will expand the study to include a larger, heterogeneous population as well as cover a greater variation of medical contexts. A larger study will enable us to create a comprehensive evaluation dataset for GenAI systems, ensuring a rigorous task-centric evaluation of GenAI systems.
The second avenue is studying a wider range of GenAI system configurations. Generative AI systems are considerably malleable and their behavior can be modulated through various prompting strategies, supervised fine tuning, and reinforcement learning etc. This paper focuses on sketching out an evaluation paradigm by studying two simple configurations. In future, we will build and study methods that align GenAI system response to experts.
@@
Third, prominent metrics such like BLEU [54], ROGUE [50], BERT Score [55], or USE [51] while commonly used to evaluate LLM responses, fall short of measuring if the response produced are human-like or if they adequately address the humans’ information needs. To address this shortcoming in AI & ML evaluation paradigm, we will develop a quantitative measurement scale based on our the qualitative evaluation paradigm proposed here. Finally, we are invested in developing intelligent assistive systems for the healthcare. Our future work will explore how the shortcoming highlighted in this paper can be alleviated with modern and classical AI methods to develop useful conversational agents.
Despite its limitations, our makes several important contributions towards sustainable and productive deployment of generative AI systems. First, the performance of generative AI systems is typically measured on domain-agnostic, task-agnostic question-answering datasets. While these methods are useful in comparing algorithms against each other, they provide little insight about how useful the answers are for humans. Our research situates question answering in a specific context of people seeking information to enhance understanding in healthcare settings. This enables us measure the degree to which GenAI systems are useful. Second, GenAI systems are typically evaluated on datasets curated by and for clinicians. These datasets contain questions and answers that are most interesting to clinicians while performance measurement includes criterion such as diagnostic accuracy. Such an approach introduces an implicit bias in the design and analysis of these systems - they are designed to replace clinicians by mimicking the output of their problem-solving process, following the AI-is-automation philosophy. However, to design a good information support tool, paying attention to who the seeker and the consumer of that information is critical. Patients and caregivers have vastly different needs than clinicians who are experts of terminology, diagnostics, and treatment. Our research takes the AI-is-augmentation view and centers patients and caregivers, adopting a much needed lens missing from the medical generative AI research landscape. Finally, this paper lays the groundwork for a principled, within-context evaluation approach for generative AI systems. We study the space of information needs of a patient/caregiver to understand what types of questions are asked by the patient/caregiver. We introduce a novel qualitative answer evaluation paradigm that measures both correctness and relevance of answers generated by AI systems. Our research can be considered a precursor to a rigorous, quantitative approach for measuring suitability of a generative AI system for addressing humans’ informational needs.